OpenStudy (cutiecomittee123):

If cosx=2/3 and x is in quadrant 4, find sin(x/2)

OpenStudy (tkhunny):

What is your plan for a solution? If x is in quadrant 4 \(270º < x < 360º\), where does that put x/2?

OpenStudy (cutiecomittee123):

not sure? halfway between 270 and 360?

OpenStudy (cutiecomittee123):

315?

OpenStudy (tkhunny):

Divide all three elements. \(270º < x < 360º\) \(135º < x/2 < 180º\) It is in the last half of the Quadrant II. Make sense?

OpenStudy (cutiecomittee123):

okay I am following so far

OpenStudy (tkhunny):

What is the nature of sine and cosine in Quadrant II? Cosine is negative and sine is positive. Don't forget this! We'll get it wrong if we do. Fair enough?

OpenStudy (cutiecomittee123):

Yes got it!

OpenStudy (tkhunny):

Now, we need a half angle formula. Have you one of those?

OpenStudy (tkhunny):

Waiting... hoping... looking it up... \(\sin(x/2) = \pm\)... Are we close?

OpenStudy (cutiecomittee123):

well I only have double angle identities in my notes which are Cos2x=1-2sin^2x, cos^2x-sin^2x, 2cos^2x-1 sin2x=2sincosx tan2x=(2tanx)/1+tan2x

OpenStudy (cutiecomittee123):

but I am not that sure about half identitites because I dont see them anywhere in my notes, just squared, double, sum, and difference identitities

OpenStudy (tkhunny):

That's good enough. Use one of those cosine ones. We need sin(x/2), so do a little algebra. \(\cos(2x) = 1 - 2\sin^{2}(x)\) Use a different angle... y = 2x to give... \(\cos(y) = 1 - 2\sin^{2}(y/2)\) -- We just invented a half-angle formula!! Still with me?

OpenStudy (cutiecomittee123):

Okay got it. I didnt know you could do that. But why the (y/2) I dont see how that came from plugging in the y=2x to the initial double angle identity for cos?

OpenStudy (cutiecomittee123):

because there is only an x there?

OpenStudy (tkhunny):

y = 2x ==> y/2 = x Just division by 2.

OpenStudy (cutiecomittee123):

Oh I see! makes sense

OpenStudy (tkhunny):

x, y, q, Fred -- It is of no consequence. It's just a label. Let's solve for ... \(\sin(y/2) = \pm\sqrt{\dfrac{1\cos(y)}{2}}\) Just algebra. Nothing tricky.

OpenStudy (cutiecomittee123):

thanks for going through this in detailed steps btw, this is exactly what I needed! this stuf is hard.

OpenStudy (cutiecomittee123):

Hmm so square both sides then multiply both sides by 2 Then divide by one and cos(y)=(sin^2 2(x))/1

OpenStudy (tkhunny):

Tough question. What does that "\(\pm\)" symbol mean?

OpenStudy (cutiecomittee123):

Well in this case we either have to chose the negative or possitive answer based on the quadrant the answer is in

OpenStudy (tkhunny):

That silly symbol means about six different things in all of mathematics. Excellent!! In this case, that \(\pm\) mean to pick the correct sign. In this case, we are in Quadrant II and we are talking about the sine, so pick \(+\). Substitute the known value of cosine and do the arithmetic.

OpenStudy (cutiecomittee123):

and in this case we found out it is in the second quadrant and we are focussing on sin so it is only the possitive

OpenStudy (cutiecomittee123):

Yes:)

OpenStudy (cutiecomittee123):

SO it would be cosy=sin^2 2x/1

OpenStudy (tkhunny):

Whoops!!! I didn't type it correctly. \(\sin(x/2) = \pm\sqrt{\dfrac{1 - \cos(x)}{2}}\). Sorry about that.

OpenStudy (cutiecomittee123):

Oh so we are using x instead?

OpenStudy (cutiecomittee123):

but then wouldnt it be the same but with x o cos(x)=(sin^2 2x)/1

OpenStudy (tkhunny):

Doesn't matter. The original problem statement gave us 'x', so we may as well go with it. Don't try to transform back to the double angle. Think of it as \(\sin(HalfofSomeAngle) = \sqrt{\dfrac{1 - \cos(SomeAngle)}{2}}\) It does not matter what you call "some angle". Just go with it. Whether it is x and x/2 or z/5 and z/10 or 12w and 6w - as long as the one outside the radical is half the size of the one inside the radical. It's the same result. \(\cos(x) = 2/3\) so \(\sin(x/2) = \sqrt{\dfrac{1 - 2/3}{2}}\) I'll leave you the arithmetic. gtg